# Data-Driven Superstabilizing Control under Quadratically-Bounded Errors-in-Variables Noise

### v1

@@ III-C Robustified Linear Program
Problem (III.1) involves $2f^{2}+f$ infinite-dimensional LP, each posed over the $n(n+T)+m(n+T-1)$ variables $(A,B,\Delta x,\Delta u)$. Reducing the number of variables involved in any quantification will simplify resultant SOS truncations and increase computational tractability.
This subsection will use a Theorem of Alternatives in order to eliminate the noise variables $(\Delta x,\Delta u)$ from Program (III.1). The maximum number of variables appearing in any infinite-dimensional linear inequality constraint will subsequently drop from $n(n+T)+m(n+T-1)$ down to $n(n+m)$. Let $q(A,B):\mathcal{P}\rightarrow\mathbb{R}$ be a function that is independent of $(\Delta x,\Delta u)$, such as any one of the left-hand-side elements for constraints (24b)-(24c). The following pair of problems are strong alternatives (exactly one is feasible):
@@
The $(A,B)$ dependence of $(\mu,s,\tau)$ will be omitted to simplify notation. A weighted sum $\Phi(A,B,\Delta x,\Delta u;\mu,s,\tau)$ may be constructed from $q$ and $(\mu,s,\tau)$ from (III-C):
@@
$\textstyle+\sum_{\ell=1}^{L}\left(s_{\ell}^{\top}(F_{\ell}\Delta x+G_{\ell}\Delta u)-\tau_{\ell}\right).$ (27)
@@
$\Phi$ $=\textstyle\left(-q+\mu^{\top}h^{0}-\sum_{\ell=1}^{L}\tau_{\ell}\right)$
$\textstyle+\left(\Xi^{\top}\mu+\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}\right)^{\top}\Delta x$ (28)
$+\textstyle\left((I_{T-1}\otimes B^{\top})\mu+\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}\right)^{\top}\Delta u.$
@@
$\sup_{\Delta x,\Delta u}\Phi=\begin{cases}Q&\textrm{if}\ \Xi^{\top}\mu+\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}\\ &\textrm{if}\ (I_{T-1}\otimes B^{\top})\mu+\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}\\ \infty&\textrm{otherwise}.\end{cases}$ (30)
Theorem III.3
@@
$\operatorname*{find}_{\mu,s,\tau}$ $\forall(A,B)\in\mathcal{P}:$ (31a)
$\qquad q+\textstyle\sum_{\ell=1}^{L}\tau_{\ell}-\mu^{\top}h^{0}\geq 0$ (31b)
$\textstyle\qquad\Xi^{\top}\mu+\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}=0$ (31c)
$\qquad\textstyle(I_{T-1}\otimes B^{\top})\mu+\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}=0$ (31d)
$(\mu,s,\tau)$ from (III-C). (31e)
@@
In the case of probabilistic noise set from (II-B1) with $L=2T-1$, the multipliers $(s,\tau)$ can be partitioned into $\forall t\in 1..T:\ (s^{x}_{t},\tau^{x}_{t})$ for (4a) and $\forall t\in 1..T-1:\ (s^{u}_{t},\tau^{u}_{t})$ for (4b). The certificate (III.3) can then be expressed as
$\operatorname*{find}_{\mu,s,\tau}$ $\forall(A,B)\in\mathcal{P}:$ (32a)
$\qquad q+\textstyle\sum_{\ell=1}^{T}\sqrt{\varepsilon(\delta_{x};n)}\tau^{x}_{\ell}+\textstyle\sum_{\ell=1}^{T-1}\sqrt{\varepsilon(\delta_{u};m)}\tau^{u}_{\ell}$
@@
$\qquad\Sigma^{-1/2}_{x}s_{1}^{x}=-A^{\top}\mu_{1}$ (32c)
$\forall t=2..T-1:\ \Sigma^{-1/2}_{x}s_{t}^{x}=\mu_{t-1}-A^{\top}\mu_{t}$ (32d)
$\qquad\Sigma^{-1/2}_{x}s_{T}^{x}=\mu_{T-1}$ (32e)
$\forall t\in 1..T-1:\ \Sigma_{u}^{-1/2}s^{u}_{t}=-B^{\top}\mu_{t}$ (32f)
$(\mu,s,\tau)$ from (III-C). (32g)
@@ IV Truncated Sum-of-Squares Programs
This section uses the moment-SOS hierarchy of SDPs to discretize the infinite-dimensional LP in (III.1) into finite-dimensional convex optimization problems that are more amenable to computation. This discretization will be performed by SOS-matrix truncations.
@@ IV-A Quadratically-Bounded Truncations
Program (III.1) has $2f^{2}+f$ infinite-dimensional LP constraints posed over $(A,B,\Delta x,\Delta u)$. This subsection applies a degree $2k$ SOS tightening to the constraints in program (III.1). In each case, the matrix $M$ is restricted to a polynomial $M\in(\mathbb{R}[A,B])^{n\times n}_{\leq 2k}.$ The remainder of this section will analyze the computational complexity of an SOS tightening on a single infinite-dimensional constraint from (III.1) (represented as $q(A,B)\geq 0$ from (25a)). Complexity will be compared according to the PSD Gram matrix of maximal size.
@@ IV-A1 Full Program
The full program applies a scalar WSOS constraint $q\in\Sigma^{1}[\bar{\mathcal{P}}]$ posed over variables $(A,B,\Delta x,\Delta u)$. The size of the maximal Gram matrix for each (25a) restriction is
@@ IV-A2 Alternatives (Dense)
$(s_{\ell},\tau_{\ell})\in\mathbb{L}^{n_{\ell}}$ $\Leftrightarrow\begin{bmatrix}\tau_{\ell}I_{n_{\ell}}&s_{\ell}\\ s_{\ell}^{\top}&\tau_{\ell}\end{bmatrix}\in\mathbb{S}^{n_{\ell}+1}_{+}.$ (36)
In order to prove convergence of the Alternative truncations as the degree $k\rightarrow\infty$, we must assume that there exists a known Archimedean set $\Pi(A,B)$ such that $\Pi\supseteq\mathcal{P}$. Such a $\Pi$ might be known from norm or Lipschitz bounds on $(A,B)$, or other similar knowledge on reasonable plant behavior. If $\Pi$ is a-priori unknown, then we will use the WSOS symbol $\Sigma^{r}[\Pi]$ to refer to the SOS set $\Sigma^{r}[x]$. At the degree-$k$ truncation, the multipliers from (III-C) can be chosen:
@@
$\begin{bmatrix}\tau_{\ell}(A,B)I_{n}&s_{\ell}(A,B)\\ s_{\ell}(A,B)^{\top}&\tau_{\ell}(A,B)\end{bmatrix}\in\Sigma^{n_{\ell}+1}[\Pi]_{\leq 2k}$ $\forall\ell\in 1..L.$ (37f)
The matrix WSOS constraint in (37f) has \@iaciPSD PSD Gram matrix of maximal size
@@
The ‘dense’ nomenclature for this approach will refer to the imposition that (37f) is WSOS over a matrix of size $n_{\ell}+1$ for each $\ell=1..L.$
@@ IV-A3 Alternatives (Sparse)
$(s_{\ell},\tau_{\ell})\in\mathbb{L}^{n_{\ell}}$ $\Leftrightarrow\exists z_{i\ell}:\ \begin{bmatrix}\tau_{\ell}&s_{i\ell}\\ s_{i\ell}&z_{i\ell}\end{bmatrix}\in\mathbb{S}^{2}_{+},\quad\tau_{\ell}=\textstyle\sum_{i=1}^{n}z_{i\ell}.$ (41)
The multipliers in (III-C) can be restricted to
@@
$z_{\ell}$ $\in(\mathbb{R}[\Pi]_{\leq 2k})^{n_{\ell}}$ $\forall\ell\in 1..L$ (42c)
$\tau_{\ell}$ $=\textstyle\sum_{i=1}^{n_{\ell}}z_{i\ell}$ $\forall\ell\in 1..L$ (42d)
@@
$\begin{bmatrix}\textstyle\tau_{\ell}(A,B)&s_{i\ell}(A,B)\\ s_{i\ell}(A,B)&z_{i\ell}(A,B)\end{bmatrix}\in\Sigma_{\leq 2k}^{2}[\Pi]$ $\forall\ell=1..L.$ (42g)
The Gram matrices from (42g) have a reduced size of
@@
The cone constraints (37f) and (42g) are not necessarily equivalent at finite degree $2k$, although they will describe the same set as $k\rightarrow\infty$ given that $(s,\tau)$ are optimization variables. Refer to [27] for more details on the relationship between matrix SOC cone representations.
@@ IV-B Computational Complexity
Computational complexity of the Full, Alternatives (Dense), and Alternatives (Decomposed) schemes will be judged comparing the sizes and multiplicities of the largest PSD constraint in any $\bar{\mathcal{P}}$-nonnegativity constraint. As a reminder, the extended superstabilizing program in (III.1) has $2n^{2}+n$ such $\bar{\mathcal{P}}$-nonnegativity linear inequality constraints.
For each $\bar{\mathcal{P}}$-nonnegativity constraint, the Full program requires 1 block of maximal size $p_{F}$ from (33). The Alternatives programs require $L+1$ PSD blocks each, in which the Dense program has block sizes of $p_{A}^{\ell}$ from (38) for $\ell=1..L$ (and a scalar block of size $\binom{n(n+m)+k}{k}$). The sparse program has all $L$ blocks of size $p_{B}$ from (43) along with a similar scalar block of size $\binom{n(n+m)+k}{k}$. Table I reports the size of the largest PSD matrix constraint for the three approaches under $n=2,m=2,k=2,T=12$. The considered quadratically-bounding constraints all have $L=2$, such as in the elementwise-norm $\lVert\Delta x_{t}\rVert_{2}\leq\epsilon_{x}$ and $\lVert\Delta u_{t}\rVert_{2}\leq\epsilon_{u}$ from Section II-B1. We first note that the per-iteration complexity of solving an SDP using an interior point method (with PSD size $N$ having $M$ affine constraints) is $O(N^{3}M+N^{2}M^{2})$ [28]. While the specifics of parameter scaling for nonsymmetric interior point SOS solvers are different (lacking the affine constraint embedding) [29], the maximal PSD size remains a valuable way to compare predicted performance. In the context of Table I, the size of $N=p_{F}=1540$ is intractably large for current interior-point methods. The ‘Count’ scaling for Dense and Sparse causes computational complexity to grow linearly, while the increasing ‘Size’ parameter causes polynomial growth in scaling.
@@
V Included Process Noise
Process noise can be added to the quadratically-bounded EIV framework developed in this paper. Consider structured process noise $\{w_{t}\}$ corrupting dynamics in (1a), leading to the relation (with known $E\in\mathbb{R}^{n\times n_{w}}$)
$x_{t+1}=Ax_{t}+Bu_{t}+Ew_{t}.$ (44)
The consistency set of plants $\bar{\mathcal{P}}^{w}(A,B,\Delta x,\Delta u,w)$ compatible with data in $\mathcal{D}$ is
$\bar{\mathcal{P}}^{w}:=\begin{Bmatrix}[l]\Xi\Delta x+(I_{T-1}\otimes B)\Delta u+(I_{T-1}\otimes E)w=-h^{0}\\ (F_{\ell}\Delta x+G_{\ell}\Delta u+H_{\ell}w,1)\in\mathbb{L}^{n_{\ell}}\qquad\forall\ell\in 1..L\\ \end{Bmatrix},$ (47)
in which $\{H_{\ell}\}$ are matrices reflecting the quadratic constraints imposed on $w$. The process-noise version of the constraint (25a) is
$\forall(A,B,\Delta x,\Delta u,w)\in\bar{\mathcal{P}}^{w}:$ $q(A,B)\geq 0.$ (48)
A degree-$2k$ SOS truncation of (48) can proceed directly by $q(A,B)\in\Sigma^{1}[\bar{\mathcal{P}}^{w}]_{\leq 2k},$ yielding Gram matrices of size $\binom{n(n+T)+m(n+T-1)+n_{w}+k}{k}$. The Theorem of Alternatives process eliminates the noise variables $(\Delta x,\Delta u,w)$ altogether. The expression of (48) adds the following entry to the constraints in (III.3):
$(I_{T-1}\otimes E^{\top})\mu+\textstyle\sum_{\ell=1}^{L}H_{\ell}^{\top}s_{\ell}=0.$ (49)
Just like in the case without process noise, the infinite-dimensional linear constraints depend solely on $(A,B)$ after robustifying against noise, and the size of the maximal Gram matrices are found in (38) and (43).

### v2

@@ III-C Robustified Linear Program
Problem (15) involves \@iaciLP LP with $2f^{2}+f$ infinite-dimensional linear constraints, each posed over the $n(n+T)+m(n+T-1)$ variables $(A,B,\Delta x,\Delta u)$. Reducing the number of variables involved in any quantification will simplify resultant SOS truncations and increase computational tractability. This subsection will use a Theorem of Alternatives in order to eliminate the noise variables $(\Delta x,\Delta u)$ from Program (15). The maximum number of variables appearing in any infinite-dimensional linear inequality constraint will subsequently drop from $n(n+T)+m(n+T-1)$ down to $n(n+m)$, which no longer depends on the number of samples $T$.
Let $q(A,B):\mathcal{P}\rightarrow\mathbb{R}$ be a function that is independent of $(\Delta x,\Delta u)$, such as any one of the left-hand-side elements for constraints (15b)-(15c). The following pair of problems are strong alternatives (exactly one is feasible):
@@
The $(A,B)$ dependence of $(\mu,s,\tau)$ will be omitted to simplify notation. A weighted sum $\Phi(A,B,\Delta x,\Delta u;\mu,s,\tau)$ may be constructed from $q$ and $(\mu,s,\tau)$ from (17):
@@
$\textstyle+\sum_{\ell=1}^{L}\left(\tau_{\ell}\ -s_{\ell}^{\top}(F_{\ell}\Delta x+G_{\ell}\Delta u)\right).$ (18)
@@
$\Phi$ $=\textstyle\left(-q+\mu^{\top}h^{0}+\sum_{\ell=1}^{L}\tau_{\ell}\right)$
$\textstyle+\left(\Xi^{\top}\mu-\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}\right)^{\top}\Delta x$ (19)
$+\textstyle\left((I_{T-1}\otimes B^{\top})\mu-\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}\right)^{\top}\Delta u.$
@@
$\sup_{\Delta x,\Delta u}\Phi=\begin{cases}Q&\textrm{if}\ 0=\Xi^{\top}\mu-\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}\\ &\textrm{if}\ 0=(I_{T-1}\otimes B^{\top})\mu-\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}\\ \infty&\textrm{otherwise}.\end{cases}$ (21)
Theorem III.2
@@
[…]
$\qquad q-\textstyle\sum_{\ell=1}^{L}\tau_{\ell}-\mu^{\top}h^{0}\geq 0$ (22b)
$\textstyle\qquad\Xi^{\top}\mu-\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}=0$ (22c)
$\qquad\textstyle(I_{T-1}\otimes B^{\top})\mu-\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}=0$ (22d)
$(\mu,s,\tau)$ from (17). (22e)
@@
Remark 2
In the case of probabilistic noise set from (4) with $L=2T-1$, the multipliers $(s,\tau)$ can be partitioned into $\forall t\in 1..T:\ (s^{x}_{t},\tau^{x}_{t})$ for (4a) and $\forall t\in 1..T-1:\ (s^{u}_{t},\tau^{u}_{t})$ for (4b). The certificate (22) can then be expressed as
[…]
$\qquad q-\textstyle\sum_{\ell=1}^{T}\sqrt{\varepsilon(\delta_{x};n)}\tau^{x}_{\ell}+\textstyle\sum_{\ell=1}^{T-1}\sqrt{\varepsilon(\delta_{u};m)}\tau^{u}_{\ell}$
@@
$\qquad\Sigma^{-1/2}_{x}s_{1}^{x}=A^{\top}\mu_{1}$ (23c)
$\forall t=2..T-1:\ \Sigma^{-1/2}_{x}s_{t}^{x}=A^{\top}\mu_{t}-\mu_{t-1}$ (23d)
$\qquad\Sigma^{-1/2}_{x}s_{T}^{x}=-\mu_{T-1}$ (23e)
$\forall t\in 1..T-1:\ \Sigma_{u}^{-1/2}s^{u}_{t}=B^{\top}\mu_{t}$ (23f)
$(\mu,s,\tau)$ from (17). (23g)
@@ IV Truncated Sum-of-Squares Programs
This section uses the moment-SOS hierarchy of SDPs to discretize the infinite-dimensional LP in (15) into finite-dimensional convex optimization problems that are more amenable to computation. This discretization will be performed by SOS-matrix truncations.
@@ IV-A Quadratically-Bounded Truncations
Program (15) has $2f^{2}+f$ infinite-dimensional LP constraints posed over $(A,B,\Delta x,\Delta u)$. This subsection applies a degree $2k$ SOS tightening to the constraints in program (15). In each case, the matrix $M$ is restricted to a polynomial $M\in(\mathbb{R}[A,B]_{\leq 2k})^{n\times n}.$ The remainder of this section will analyze the computational complexity of an SOS tightening on a single infinite-dimensional constraint from (15) (represented as $q(A,B)\geq 0$ from (16a)). Complexity will be compared according to the PSD Gram matrix of maximal size.
@@ IV-A1 Full Program
The full program applies a scalar WSOS constraint $q\in\Sigma^{1}[\bar{\mathcal{P}}]$ over $(A,B,\Delta x,\Delta u)$. The size of the maximal Gram matrix for each (16a) restriction is
@@ IV-A2 Alternatives (Dense)
$(s_{\ell},\tau_{\ell})\in\mathbb{L}^{n_{\ell}}$ $\Leftrightarrow\begin{bmatrix}\tau_{\ell}&s_{\ell}^{\top}\\ s_{\ell}&\tau_{\ell}I_{n_{\ell}}\end{bmatrix}\in\mathbb{S}^{n_{\ell}+1}_{+}.$ (25)
In order to prove convergence of the Alternative truncations as the degree $k\rightarrow\infty$, we must assume that there exists a known Archimedean set $\Pi(A,B)$ such that $\Pi\supseteq\mathcal{P}$. Such a $\Pi$ might be known from norm or Lipschitz bounds on $(A,B)$, or other similar knowledge on reasonable plant behavior. If $\Pi$ is a-priori unknown, then we will use the WSOS symbol $\Sigma^{r}[\Pi]$ to refer to the SOS set $\Sigma^{r}[x]$. At the degree-$k$ truncation, the multipliers from (17) can be chosen:
@@
$\begin{bmatrix}\tau_{\ell}(A,B)&s_{\ell}(A,B)^{\top}\\ s_{\ell}(A,B)&\tau_{\ell}(A,B)I_{n}\end{bmatrix}\in\Sigma^{n_{\ell}+1}[\Pi]_{\leq 2k}$ $\forall\ell\in 1..L.$ (26d)
The matrix WSOS constraint in (26d) has \@iaciPSD PSD Gram matrix of maximal size
@@
The ‘dense’ nomenclature for this approach will refer to the imposition that (26d) is WSOS over a matrix of size $n_{\ell}+1$ for each $\ell=1..L.$
@@ IV-A3 Alternatives (Sparse)
$(s_{\ell},\tau_{\ell})\in\mathbb{L}^{n_{\ell}}$ $\Leftrightarrow\exists z_{i\ell}:\ \begin{bmatrix}\tau_{\ell}&s_{i\ell}\\ s_{i\ell}&z_{i\ell}\end{bmatrix}\in\mathbb{S}^{2}_{+},\quad\tau_{\ell}=\textstyle\sum_{i=1}^{n}z_{i\ell}.$
The multipliers in (17) can be restricted to
@@
$s_{\ell},\ z_{\ell}$ $\in(\mathbb{R}[\Pi]_{\leq 2k})^{n_{\ell}}$ $\forall\ell\in 1..L$ (28b)
@@
$\begin{bmatrix}\textstyle\sum_{i=1}^{n_{\ell}}z_{i\ell}(A,B)&s_{i\ell}(A,B)\\ s_{i\ell}(A,B)&z_{i\ell}(A,B)\end{bmatrix}\in\Sigma^{2}[\Pi]_{\leq 2k}$ $\forall\ell=1..L.$ (28c)
The Gram matrices from (28c) have a reduced size of
@@
The cone constraints (26d) and (28c) are not necessarily equivalent at finite degree $2k$, although they will describe the same set as $k\rightarrow\infty$ given that $(s,\tau)$ are optimization variables. Refer to [28] for more details on the relationship between matrix SOC cone representations.
@@ IV-B Computational Complexity
Computational complexity of the Full, Alternatives (Dense), and Alternatives (Decomposed) schemes will be judged comparing the sizes and multiplicities of the largest PSD constraint in any $\bar{\mathcal{P}}$-nonnegativity constraint. As a reminder, the superstabilizing program in (15) has $2n^{2}+n$ such $\bar{\mathcal{P}}$-nonnegativity linear inequality constraints.
For each $\bar{\mathcal{P}}$-nonnegativity constraint, the Full program requires 1 block of maximal size $p_{F}$ from (24). The Alternatives programs require $L+1$ PSD blocks each, in which the Dense program has block sizes of $p_{A}^{\ell}$ from (27) for $\ell=1..L$ (and a scalar block of size $\binom{n(n+m)+k}{k}$). The sparse program has all $L$ blocks of size $p_{B}$ from (29) along with a similar scalar block of size $\binom{n(n+m)+k}{k}$. Table I reports the size of the largest PSD matrix constraint for the three approaches under $n=2,m=2,k=2,T=12$. The considered quadratically-bounding constraints all have $L=2$, such as in the elementwise-norm $\lVert\Delta x_{t}\rVert_{2}\leq\epsilon_{x}$ and $\lVert\Delta u_{t}\rVert_{2}\leq\epsilon_{u}$ from Section II-B1. We first note that the per-iteration complexity of solving an SDP using an interior point method (with PSD size $N$ having $M$ affine constraints) is $O(N^{3}M+N^{2}M^{2})$ [29]. In the context of Table I, the size of $N=p_{F}=1540$ is intractably large for current interior-point methods. The ‘Multiplicity’ scaling for Dense and Sparse causes computational complexity to grow linearly, while the increasing ‘Size’ parameter causes polynomial growth in scaling.
@@
V Extended Superstability
The superstabilization method considered in Sections III and IV rely on the use of a previously given and fixed $W$ matrix. The framework of Extended Superstability [20] allows for $W$ to be chosen as a positive-definite $n$-dimensional diagonal matrix, in which this diagonal-restricted $W$ may be searched over in optimization and is not fixed in advance. The resultant common Lyapunov function $\lVert Wx\rVert_{\infty}$ therefore has hyper-rectangular sublevel sets. Letting $v\in\mathbb{R}^{n}$, $v>0$ be a positive vector with matrix $W=\textrm{diag}(v)^{-1},$ and $x_{t+1}=Ax_{t}$ be a dynamical system, the $W$-superstabilization condition $\lVert WAW^{+}\rVert_{\infty}<1$ may be expressed as
$\forall i=1..n,j=1..n$ $[A\ \textrm{diag}(v)]_{ij}\leq M_{ij}$ (30a)
$\forall i=1..n$ $\textstyle\sum_{j=1}^{n}M_{ij}<v_{i}.$ (30b)
The quadratically-bounded EIV extended superstabilization task involves the following optimization problem with variables $M(A,B):\mathcal{P}\rightarrow\mathbb{R}^{n\times n},S\in\mathbb{R}^{m\times n},v\in\mathbb{R}^{n}:$
[…]
$\quad\forall i\in 1..n:$ (31b)
$\quad\textstyle\qquad v_{i}>0,\ \sum_{j=1}^{n}M_{ij}(A,B)<v_{i}$
$\quad\forall i\in 1..n,\ j\in 1..n:$ (31c)
$\quad\qquad\textstyle\lvert A_{ij}v_{j}+\sum_{\ell=1}^{m}B_{i\ell}S_{\ell j}\rvert\leq M_{ij}(A,B)$
$M:\mathcal{P}\rightarrow\mathbb{R}^{n\times n},\,S\in\mathbb{R}^{m\times n},v\in\mathbb{R}^{n}.$ (31d)
If Program (31) is feasible, then the controller $K$ recovered by $K=SW=S(\textrm{diag}(v))^{-1}$ is guaranteed to $W$-superstabilize all systems in $\mathcal{P}$. All three methods (Full, Dense, Sparse) for WSOS truncation can be employed to formulate order-$k$ versions of Program (31) (w.r.t. a Psatz in $(A,B,\Delta x,\Delta u)$ for Full or $(A,B)$ for Alternatives) by choosing $M\in(\mathbb{R}[A,B]_{\leq 2k})^{n\times n}$.